# SlowFast-SCI: Slow-Fast Deep Unfolding Learning for Spectral Compressive Imaging

### v1

@@ 4.4 Ablation Studies
Fast Learning Ablation. We analyze the performance of FAST-TTAM and TTA in the fast learning. Tab. 3 (c, d) and (f, g) show that the proposed FAST-TTAM can effectively transfer the distilled model to the target domain through self-supervised training. Besides, case (g, h) in Tab. 3 shows self-supervised test-time adaptation can further align the fast unfolding variant with target scenes.
@@ 5 Conclusion
We have introduced SlowFast-SCI, to our knowledge the first test-time adaptation–driven, self-adaptive spectral unfolding framework. By adopting a dual-speed paradigm—distilling a priors-driven backbone into a compact “fast unfolding” network during offline training, then embedding lightweight, self-supervised adapters for on-the-fly calibration—our method achieves over 70% parameter and FLOPs reduction, up to 5.79 dB PSNR improvement on out-of-distribution data, and a 4× faster adaptation runtime, all while preserving robust cross-domain performance.
Crucially, SlowFast-SCI’s modularity allows it to be seamlessly integrated into any deep unfolding architecture, unlocking real-time hyperspectral reconstruction on resource-constrained platforms—ranging from aerial drones and point-of-care scanners to distributed environmental sensors. We believe this work paves the way for a new class of self-adaptive, field-deployable imaging systems and look forward to exploring extensions to other computational imaging modalities and richer adaptation objectives.
Limitation and Future Work: SlowFast-SCI’s per-sample adapters, despite their small footprint, add inference overhead that may hinder use on ultra–low-power or hard real-time platforms. Our study is limited to disperser-based CASSI; applying and validating the dual-speed approach on other SCI architectures and unfolding models remains crucial. Future directions include hardware-aware adapter designs, early-exit adaptation schemes, and extending the framework to snapshot multispectral, fluorescence, tomographic imaging and general image restoration.

### v2

@@ 6 Conclusion
We have introduced SlowFast-SCI, to the best of our knowledge the first test-time adaptation–driven, self-adaptive spectral unfolding framework. By adopting a dual-speed paradigm—distilling a priors-driven backbone into a compact “fast unfolding” network during offline training, then embedding lightweight, self-supervised adapters for on-the-fly calibration—our method achieves over 70% parameter and FLOPs reduction, up to 5.79 dB PSNR improvement on out-of-distribution data, and a 4× faster adaptation runtime, all while preserving robust cross-domain performance.
Crucially, SlowFast-SCI’s modularity allows it to be seamlessly integrated into any deep unfolding architecture, unlocking real-time hyperspectral reconstruction on resource-constrained platforms—ranging from aerial drones, point-of-care scanners to distributed environmental sensors. We believe this work paves the way for a new class of self-adaptive, field-deployable imaging systems and look forward to exploring extensions to other computational imaging modalities and richer adaptation objectives.
Limitation and Future Work: Despite their small footprint, SlowFast-SCI’s per-sample adapters introduce minor latency, posing challenges for ultra–low-power or real-time deployment. Our study focuses on disperser-based CASSI; future work will extend the dual-speed paradigm to other SCI forms and unfolding models, with emphasis on hardware-efficient adapters, early-exit schemes, and broader imaging applications.